# MicroSSIM: Improved Structural Similarity for Comparing Microscopy Data

### v1

@@ Abstract
Figure 1: Failure mode of SSIM on Microscopy Data. (Top) A noisy microscopy image, i.e. a micrograph, its denoised version predicted using N2V, and the corresponding High-SNR (noise free) ground truth is shown. There is a problem in the evaluation of denoising quality, which is that the pixel intensity distribution of the ground truth and the prediction (as shown in respective insets) are very different. This is specifically true for the foreground content which comprises brighter pixels. So applying SSIM directly on it will not give a sensible value. We solve this issue with $\mathbb{M}\text{icroSSIM}$. (Bottom) We show one example to demonstrate an apparent counter-intuitive behavior of SSIM. The SSIM between a natural image (taken from Imagenet [4]) and a pure noisy image drawn from the uniform distribution is much lower than the SSIM between a micrograph and a noisy image with identical distribution as before. The expectation naturally is to have $\text{SSIM}\approx 0$ in both cases. We solve this issue with $\mathbb{M}\text{icroSSIM}$ and appropriate data pre-processing and show it in bottom right plot where 30 random microscopy images are used.
@@ 1 Introduction
Finally, our implementation is such that it is easily extendable to other SSIM variants. To this end, we extend MS-SSIM to $\mathbb{M}\text{icroMS}3\text{IM}$. We provide the code with this manuscript which we will convert into a Python package.
@@ 2 Related work
The origin of the Structural Similarity measure (SSIM) could be traced back to the universal quality index (UQI) [11]. In this seminal work, the image similarity was decomposed into three measurable quantities namely luminance($l$), contrast ($c$), and structure($st$). Later, constants $c_{1},c_{2},c_{3}$ were added to each of the three measurable quantities proposed in UQI formulation to avoid division by zero [12]. When computing SSIM between two images $x$ and $y$, individual terms had the following expression.
@@
where $u_{z}=\sum_{k}^{K}\frac{z[k]}{K}$, $s_{z}=\sqrt{\sum_{k}^{K}\frac{(z[k]-\mu_{z})^{2}}{K-1}}$ and $s_{x,y}=\frac{\sum_{k}^{K}(x[k]-\mu_{k})(y[k]-\mu_{k})}{K-1}$. For mathematical convenience, $c_{3}=c_{2}/2,\alpha=\beta=\gamma=1.0$ was introduced which enabled simplification of the expression and it is this formulation which people refer to as SSIM today. Mathematically, the SSIM measure between two images $x$ and $y$ is computed as follows:
@@
Typically SSIM is computed over small regions and not over the entire image. To get a single value for the entire image MSSIM is used which is the average of all SSIM values computed over all small regions of the image. In practice, the MSSIM measure is what researchers refer to even when they use the term SSIM. Technically speaking,
@@ 3.1 Saturation
$g=\frac{a+c_{i}}{b+c_{i}},$ (4)
where $\{c_{i},i\in\{1,2,3\}\}$, as defined at Sec. 2, were introduced for stability. They are in turn computed from the $\gamma$ and a set of hyper-parameters $k_{i},i\in\{1,2,3\}$ as $c_{i}=(k_{i}\times\gamma)^{2}$. Now, as $\gamma$ becomes large, so does $c_{i}$. The problem arises when $a,b$ do not, for some reason, become equally large. To drive home the point, one could observe the following extreme situation: given $a,b\in R$, $\lim_{c_{i}\to\infty}g=1$, irrespective of what $a,b$ are, i.e., irrespective of how similar the two images are. We coin the term saturation to denote this aspect of SSIM components where the value becomes less sensitive to the (dis)similarity between the two images. We design the following expression to quantify it:
@@ 3.2 Handling background and large pixel intensities
To handle the offset issue, we do a pre-processing step and do background subtraction. We compute the background offset $\beta$ separately for the GT and prediction. Micrographs often have the lowest pixel value well below the average background pixel value, which is also the case in the dataset we work with. Using the minimum value would therefore be an incorrect estimate for background. So, we assume $3^{rd}$ percentile of the pixel intensities to be the background. We compute the percentile not on individual images, but rather on the whole dataset. Since background regions typically occupy more than three percent of all pixels in a micrograph, we find our empirical choice reasonable since the value changes slowly around this percentile. For instance, $0^{th}$, $1^{st}$, $3^{rd}$ and $9^{th}$ percentile for Actin GT micrographs used in this study are 41, 100, 102, and 105 respectively. However, for more densely populated micrographs, a lower percentile might be optimal. Alternatively, one can manually inspect a few micrographs and estimate an average pixel value for the background. While all of these are reasonable options, what is important to note is that when comparing multiple predictors on any dataset, the same offset must be used for GT. This is because the SSIM score depends on the offset, which we show in Supp. Sec. S1.
@@ 3.3 Optimization to obtain the scaling factor $\alpha$
Moreover, in Supp. Sec. S2, we empirically show over multiple denoising tasks that a unique maxima exists in the general case as well.
@@
It is important to stress the fact that we learn a single scalar for the entire dataset. Had we optimized for every $(x,y)$ pair, we would get a higher measure value on average, but this does not align well with the motivation for this measure, which is to estimate an optimal linear transformation between the space of predictions to their corresponding high-SNR micrographs. Naturally, this linear transformation should be the same for every GT image and prediction pair in the dataset.
@@
4 Dataset and training details
We have selected high-SNR and corresponding low-SNR images from two channels, namely Actin and Mitochondria, from the publicly available Hagen et al [5]. For training $\text{denoi}\mathbb{S}\text{plit}$, we use the same hyper-parameters along with the same train/val/test split as proposed in [1]. For $\text{denoi}\mathbb{S}\text{plit}$, we evaluate on the test set. For N2V, we evaluate on all 100 frames.
@@ 5 Experiments
Figure 5: Sensitivity to Noise on Mitochondria denoising: SSIM and $\mathbb{M}\text{icroSSIM}$ are computed between GT(col 6) and $\text{pred}_{w}$, where $\text{pred}_{w}=\text{pred}*w+\text{noise}*(1-w)$ and pred is the denoised prediction by N2V. $\text{pred}_{w}$ is shown in col 2-5 and input to N2V in col 1. Smaller $w$ means more noise in $\text{pred}_{w}$. While SSIM value barely changes across different $w$, $\mathbb{M}\text{icroSSIM}$ is more sensitive to $w$, i.e., to prediction quality.
@@
Sensitivity to noise
In order for $\mathbb{M}\text{icroSSIM}$ to be a useful measure, we need to show its practical utility. In this experiment, we report $\mathbb{M}\text{icroSSIM}$ values computed between ground truth and denoised prediction of varying quality. We show that $\mathbb{M}\text{icroSSIM}$ is more sensitive to prediction quality as compared to SSIM. For this experiment, we first use the N2V to get the denoised prediction, referred to henceforth as pred. We then mix in pred a pixel-wise independent pure noisy image, $I_{n}$, sampled from the uniform distribution by different amounts. Specifically, using a scalar $w$, we compute the inferior denoised prediction as $\text{pred}_{w}=\text{pred}*w+I_{n}*(1-w)$. A linear transformation has been applied on $I_{n}$ so as to match the predicted denoised images’ mean and standard deviation. This ensures that the mean of $\text{pred}_{w}$ does not change with $w$. This procedure produces an ordered sequence of inferior denoised predictions. In Fig. 5 and Supp. Fig. S2 we report SSIM and $\mathbb{M}\text{icroSSIM}$ on a few randomly chosen crops. While the SSIM value barely changes as we evaluate more noisy predictions, we see that $\mathbb{M}\text{icroSSIM}$ has a considerably larger reduction in score. Next, we trace the cause of this in-sensitivity of SSIM to the phenomenon of Saturation.
@@
In Supp. Sec. S1, we show that the offset value, set by the microscopist, changes the SSIM score. This is not desirable because we want the SSIM score to remain unchanged regardless of the offset, as it can be arbitrarily set by the microscopist. This experiment justifies our pre-processing step of background removal. In Supp. Sec. S2, we empirically show the uniqueness of $\alpha$. This experiment allowed us to use an off-the-shelf optimizer for the estimation of $\alpha$.
@@ Acknowledgements
Figure S1: (Left: Role of offset in luminance) In this ablation, after doing our proposed normalization of GT and prediction, we add an offset to both GT and prediction and plot the mean SSIM and mean of its components. As can be inferred from the two plots made using two random full frames from Actin (row 1) and Mito (row 2) subdatasets, offset indeed influences the luminance, and therefore, the M-SSIM value. This means that the detector offset set by the microscopist will play a role in SSIM computation. Ideally, only denoising model’s performance should play a role. We ensure this by removing the background in our normalization. (Right:Uniqueness of $\alpha$) Here, we took 30 random full frames from Actin and Mito subdataset. Using the pre-trained N2V model, we obtained the denoised predictions and used MicroSSIM to evaluate the prediction. Instead of estimating the $\alpha$, we manually tried different $\alpha$s and computed MicroSSIM using it. One can see that for every curve, a unique $\alpha$ exists which gives the maximal SSIM. This uniqueness allows us to use any off-the-shelf optimizer to obtain this maximal $\alpha$.
Figure S2: Sensitivity to Noise on Actin denoising: SSIM and $\mathbb{M}\text{icroSSIM}$ are computed between GT(last column) and $\text{pred}_{w}$, where $\text{pred}_{w}=\text{pred}*w+\text{noise}*(1-w)$. pred is the denoised prediction by N2V. $\text{pred}_{w}$ is shown in columns 2-5. Input to N2V is shown in column 1. Smaller $w$ means more noise in $\text{pred}_{w}$. SSIM value barely changes across different $w$. Compared to SSIM, our $\mathbb{M}\text{icroSSIM}$ is more sensitive to noise.
S1 Inspecting role of offset ($\beta$) on vanilla SSIM
We model the offset added by the microscope with a single scalar value added to both GT and the prediction. In Fig. S1, we investigate the case when the same offset is present in both high-SNR and low-SNR images. For this experiment, we randomly pick a GT and the corresponding N2V’s prediction pair and normalize them as is done for $\mathbb{M}\text{icroSSIM}$. We then create multiple (prediction, GT) pairs by adding a specific offset to the original GT and prediction. We compute the SSIM and its components for every offset and plot the curves. As can be seen from Fig. S1, the offset changes the luminance, and therefore the SSIM value. This observation is important because this means that the SSIM value, and therefore, the quantification of a denoiser’s performance, depends upon the offset added by the detector. This is certainly undesirable. Our pre-processing operation of background subtraction removes this offset and therefore removes this unwanted dependence on the SSIM value. While it is common to normalize by removing the mean, from Fig. 2S1, we show that removing the background is a better option for microscopy data.
S2 Inspecting uniqueness of $\alpha$
In Fig. S1, we do a grid search and vary the $\alpha$ from 00 to $300$. As can be seen from the plot, a unique scaling factor exists that yields the optimal SSIM value. This observation enabled us to use an off-the-shelf optimizer to estimate this scaling factor.
S3 Potential limitations
@@
In Fig. 3 and Supp. Fig. S4, the performance of the baseline Inst-bkg-rm raises questions about the optimization procedure. Inst-bkg-rm attains a high SSIM value between a micrograph and a pure noisy image. What this means is that if the pixel count over which the optimization is done is relatively less, then such an $\alpha$ could be obtained which will make $\mathbb{M}\text{icroSSIM}$ produce a high score when operated on dissimilar image pairs.
S4 Derivation for optimality of $\alpha$

### v2

@@ Abstract
Figure 1: Failure mode of SSIM on Microscopy Data. (A) A noisy microscopy image, i.e. a micrograph, its denoised version predicted using N2V, and the corresponding High-SNR (noise free) ground truth is shown. There is a problem in the evaluation of denoising quality, which is that the pixel intensity distribution of the ground truth and the prediction (as shown in respective insets) are very different. This is specifically true for the foreground content which comprises brighter pixels. So applying SSIM directly on it will not give a sensible value. We solve this issue with $\mathbb{M}\text{icroSSIM}$. (B) We show one example to demonstrate an apparent counter-intuitive behavior of SSIM. The SSIM between a natural image (taken from Imagenet [4]) and a pure noisy image drawn from the uniform distribution is much lower than the SSIM between a micrograph and a noisy image with identical distribution as before. The expectation naturally is to have $\text{SSIM}\approx 0$ in both cases. We solve this issue with $\mathbb{M}\text{icroSSIM}$ and appropriate data pre-processing and show it in bottom right plot where 30 random microscopy images are used.
@@ 1 Introduction
Finally, our implementation is such that it is easily extendable to other SSIM variants. To this end, we extend Multiscale SSIM (MS-SSIM) [10] to $\mathbb{M}\text{icroMS}3\text{IM}$.
@@ 2 Related work
The origin of the Structural Similarity measure (SSIM) could be traced back to the universal quality index (UQI) [11]. In this seminal work, the image similarity was decomposed into three measurable quantities namely luminance($l$), contrast ($c$), and structure($st$). Later, constants $c_{1},c_{2},c_{3}$, which depended on the range of pixel values, were added to each of the three measurable quantities proposed in UQI formulation to avoid division by zero [12]. When computing SSIM between two images $x$ and $y$, individual terms had the following expression.
@@
where $u_{z}=\sum_{k}^{K}\frac{z[k]}{K}$, $s_{z}=\sqrt{\sum_{k}^{K}\frac{(z[k]-\mu_{z})^{2}}{K-1}}$ and $s_{x,y}=\frac{\sum_{k}^{K}(x[k]-\mu_{k})(y[k]-\mu_{k})}{K-1}$. Here, $z\in\{x,y\}$. For mathematical convenience, $c_{3}=c_{2}/2,\alpha=\beta=\gamma=1.0$ was introduced which enabled simplification of the expression and it is this formulation which people refer to as SSIM today. Mathematically, the SSIM measure between two images $x$ and $y$ is computed as follows:
@@
Typically, SSIM is not computed over the entire image but rather over small overlapping regions in an efficient way using convolution operation. The result is that one has an SSIM value for every pixel of the predicted image which is sufficiently away from the image boundary. To get a single value for the entire image, MSSIM is used which is the average of all SSIM values computed over all small regions of the image. In practice, the MSSIM measure is what researchers refer to even when they use the term SSIM. Technically speaking,
@@ 3.1 Saturation
$g_{i}=\frac{a+c_{i}}{b+c_{i}},$ (4)
where $\{c_{i},i\in\{1,2,3\}\}$, as defined at Sec. 2, were introduced for stability. All $c_{i}$ values are in turn computed from the $\gamma$ and a set of hyper-parameters $k_{i},i\in\{1,2,3\}$ as $c_{i}=(k_{i}\times\gamma)^{2}$. Now, as $\gamma$ becomes large, so does $c_{i}$. The problem arises when $a,b$ do not, for some reason, become equally large. To drive home the point, one could observe the following extreme situation: given $a,b\in R$, $\lim_{c_{i}\to\infty}g=1$, irrespective of what $a,b$ are, i.e., irrespective of how similar the two images are. We coin the term saturation to denote this aspect of SSIM components where the value becomes less sensitive to the (dis)similarity between the two images. We design the following expression to quantify it:
@@ 3.2 Handling background and large pixel intensities
To handle the offset issue, we do a pre-processing step and do background subtraction. We compute the background offset $\beta$ separately for the GT and prediction. Micrographs often have the lowest pixel value well below the average background pixel value. Using the minimum value would therefore be an incorrect estimate for background. So, we assume $3^{rd}$ percentile of the pixel intensities to be the background. We compute the percentile not on individual images, but rather on the whole dataset. Since background regions typically occupy more than three percent of all pixels in a micrograph, we find our empirical choice reasonable since the value changes slowly around this percentile. For instance, $0^{th}$, $1^{st}$, $3^{rd}$ and $9^{th}$ percentile for Actin GT micrographs used in this study are 41, 100, 102, and 105 respectively. However, for more densely populated micrographs, a lower percentile might be optimal. Alternatively, one can manually inspect a few micrographs and estimate an average pixel value for the background. While all of these are reasonable options, what is important to note is that when comparing multiple predictors on any dataset, the same offset must be used for GT. This is because the SSIM score depends on the offset, which we show in Sec. 4.
@@ 3.3 Optimization to obtain the scaling factor $\alpha$
Moreover, in Sec. 4, we empirically show over multiple denoising tasks that a unique maxima exists in the general case as well.
@@
Here, $max_{GT}$ is the maximum pixel intensity observed in all GT images. It is important to stress the fact that we learn a single scalar for the entire dataset. Had we optimized for every $(x,y)$ pair, we would get a higher measure value on average, but this does not align well with the motivation for this measure, which is to estimate an optimal linear transformation between the space of predictions to their corresponding high-SNR micrographs. Naturally, this linear transformation should be the same for every GT image and prediction pair in the dataset.
@@ 4 Experiments
Dataset and training details
We selected high-SNR and corresponding low-SNR images of Actin and Mitochondria channels taken from the publicly available Hagen et al [5]. For training $\text{denoi}\mathbb{S}\text{plit}$, we use the same code, hyper-parameters, and train/val/test split as proposed in [1]. For $\text{denoi}\mathbb{S}\text{plit}$, we evaluate the test set. For N2V, we evaluate on all 100 frames. We used CAREamics library  [^1] https://github.com/CAREamics for training N2V.
@@
Figure 5: (Left: Role of offset in luminance) In this ablation, after doing our proposed normalization of GT and prediction, we add an offset to both GT and prediction and plot the mean SSIM and mean of its components. As can be inferred from the two plots made using two random full frames from Actin (row 1) and Mito (row 2) subdatasets, offset indeed influences the luminance, and therefore, the M-SSIM value. This means that the detector offset set by the microscopist will play a role in SSIM computation. Ideally, only denoising model’s performance should play a role. We ensure this by removing the background in our normalization. (Right:Uniqueness of $\alpha$) Here, we took 30 random full frames from Actin and Mito subdataset. Using the pre-trained N2V model, we obtained the denoised predictions and used MicroSSIM to evaluate the prediction. Instead of estimating a single value of $\alpha$ for all images, we manually tried different $\alpha$ values and computed SSIM using it. One can see that for every curve, a unique $\alpha$ exists which gives the maximal SSIM.
@@
Inspecting role of offset ($\beta$) on vanilla SSIM
We model the offset added by the microscope with a single scalar value added to both GT and the prediction. In Fig. 5, we investigate the case when the same offset is present in both high-SNR and low-SNR images. For this experiment, we randomly pick a GT and the corresponding N2V’s prediction pair and normalize them as is done for $\mathbb{M}\text{icroSSIM}$. We then create multiple (prediction, GT) pairs by adding a specific offset to the original GT and prediction. We compute the SSIM and its components for every offset and plot the curves. As can be seen from Fig. 5, the offset changes the luminance, and therefore the SSIM value. This observation is important because this means that the SSIM value, and therefore, the quantification of a denoiser’s performance, depends upon the offset added by the detector. This is certainly undesirable. Our pre-processing operation of background subtraction removes this offset and therefore removes this unwanted dependence on the SSIM value. While it is common to normalize by removing the mean, results in Fig. 4 and (CARE-SSIM results) in Fig. 2 highlights issues with mean-removal.
Inspecting uniqueness of $\alpha$
In Fig. 5 (right), we do a grid search and vary the $\alpha$ from 00 to $300$. As can be seen from the plot, a unique scaling factor exists that yields the optimal SSIM value. This observation enabled us to use an off-the-shelf optimizer to estimate this scaling factor.
@@ Acknowledgements
S1 Limitations and future work
@@
We argue that saturation is the critical reason that causes high SSIM score between a micrograph and an image containing pure noise. While we showed that $\mathbb{M}\text{icroSSIM}$ has lower saturation than other SSIM variants, $\mathbb{M}\text{icroSSIM}$ does not explicitly has any formulation to control saturation. So, one avenue for improvement is to develop better optimization procedures that can more effectively limit saturation. A different direction can be to find systematic ways to obtain the constants $c_{1},c_{2}$, and $c_{3}$ with the objective of having low saturation.
S2 Derivation for optimality of $\alpha$